# Simple Yet Effective Selective Imputation for Incomplete Multi-view Clustering

### v1

@@ III-B Deep Multi-view Gaussian Mixture Model
In the generative process described above, the cluster assignment $\bm{c}_{i}$ first determines the latent representation $\bm{z}_{i}$, which serves as the underlying variable that generates multi-view observations $\bm{x}_{i}^{v}$}.
During inference, the direction is reversed: we aim to infer $\bm{z}_{i}$ and $\bm{c}_{i}$ from the observed views $\bm{x}_{i}^{v}$}. Specifically, we first infer a latent representation $\bm{z}_{i}$ by integrating information from all available views, and then estimate the cluster assignment $\bm{c}_{i}$ based on the inferred $\bm{z}_{i}$. Hence, the learning objective focuses on obtaining accurate and stable latent representations rather than reconstructing all missing observations. However, incomplete views provide limited evidence, which leads to uncertain or biased estimates of posterior distribution of $\bm{z}_{i}$ and consequently degrades clustering performance.
There are two intuitive strategies to handle missing data. One is to directly recover missing data from the latent representation $\bm{z}_{i}$. However, such distribution-driven imputation contributes little to learning a more discriminative and cluster-consistent $\bm{z}_{i}$, and thus brings limited benefit to our clustering objective. Another strategy is to ignore the missing views and only aggregate observed ones. Yet under high missing rates or severe imbalance, this approach suffers from drastic degradation, and in extreme cases degenerates into a single-view solution that still outputs over-confident but biased representations.
@@
To address this, we introduce two complementary components: informativeness-based imputation position selection and distribution-level imputation. The former identifies positions with abundant and trustworthy evidence, ensuring that only well-supported missing positions are imputed, thereby mitigating the negative impact of unreliable auxiliary information. The latter models the uncertainty inherent in the imputation process by aggregating information from local neighbors at the distributional level, which not only stabilizes the missing-view likelihood approximation but also facilitates the learning of more robust and cluster-consistent latent representations.
III-C Informativeness-Based Imputation Position Selection
@@
Building upon these two complementary evidence sources, we construct a unified estimation framework that quantifies their information contribution via a position-specific support set. This allows us to compute an information score $\mathrm{Info}(i,v)$ for each missing position and perform imputation only when sufficient support is available.
This quantification is inspired by the generative assumption of DMGMM, where samples with more similar latent posteriors are more likely to belong to the same Gaussian component (i.e., cluster). From this perspective, the contribution of different reference samples and views to imputing a missing position $\tilde{\bm{x}}_{i}^{v}$ can be naturally distinguished:
@@
We quantify the informativeness of missing positions via the above principles. Each sample-view pair contributes differently based on its availability, similarity, and alignment with the target view. Specifically, the informativeness-based imputation position selection includes defining support samples, estimating information contributions, and computing information scores. Next, each part will be introduced in detail.
@@
In practice, not all samples are suitable for participating in the quantification of informativeness for a missing position. As illustrated on the right side of Fig. 3, samples that exhibit completely different missing patterns or lack any overlap in observed views should be excluded. This exclusion is not due to unreliability of the data, but rather because these samples do not provide a clear pathway for estimating the target view.
In most imputation paradigms, the estimation of a missing position relies on observable correspondences: either intra-view neighbors that share similar features, or cross-view relations established through jointly observed samples. When two samples share no common observed view, their potential statistical dependencies can only be implicitly captured during global network training, which is indirect, unstable, and not interpretable at the instance level. Therefore, we construct a set of support samples that serve as valid reference points for each missing position $(i,v)$.
@@
1) View correlation. To quantify the consistency of shared semantics across views and provide reliable weighting for similarity approximation when the target view is missing, the inter-view correlation $\mathrm{corr}^{uv}\in(0,1]$ is computed in the latent space via Canonical Correlation Analysis (CCA) [andrewDeep2013]. Specifically, for views $u$ and $v$, we consider the $N_{uv}=|\mathcal{X}^{u}\cap\mathcal{X}^{v}|$ samples observed in both views and form their latent representations:
@@
where $\bm{\mu}^{v}$ denotes the mean vector output by the encoder $f^{v}(\bm{x}^{v})$. The pre-trained encoder thus provides a well-initialized latent representation space that facilitates both reliable informativeness estimation and faster convergence in the following learning stage. Note that while using CCA, our framework is agnostic to the specific choice of correlation measure. For different models, better options such as mutual information or cross-view mapping loss could be preferred.
2) Sample similarity. To differentiate the contributions of intra-view neighbors and prioritize trustworthy samples more likely to belong to the same cluster or Gaussian component, intra-view similarity is defined as a normalized distance metric. Specifically, intra-view similarity $\mathrm{sim}_{ij}^{u}$ is computed as:
@@
Information Score Computation
@@
For cross-view evidence $\bm{x}_{i}^{u}$ ($u\neq v$), the effectiveness of these evidence hinges on the quality of view-to-view semantic transfer, such as how well the mapping between views $u$ and $v$ has been learned. As this direct quality is challenging to measure, we approximate the contribution of $\bm{x}_{i}^{u}$ indirectly by aggregating information score from co-occurring support samples weighted by view correlation and local similarity:
@@
By setting $\mathrm{corr}^{vv}=1$, the total information score for $\tilde{\bm{x}}_{i}^{v}$ is the sum of intra- and cross-view score:
@@
We use this scalar value to selectively impute only those missing positions that are sufficiently supported. In detail, we determine the threshold $\tau$ connected with a proportion of missing data, and only perform imputation when $\mathrm{Info}(i,v)>\tau$.
III-D Variational Inference with Partial Imputation
@@
Common Representation
@@
where $\bm{\mu}_{i}$ and $(\bm{\sigma}_{i})^{2}$ denote the aggregated mean and variance across available views $\mathcal{V}_{i}$. This weights views by confidence (lower variance), but high missing rates or unbalanced view missing can destabilize the posterior. Missing views reduce terms in Eq. (15), skewing $\bm{\mu}_{i}$ and underestimating $(\bm{\sigma}_{i})^{2}$, especially when views are systematically absent, leading to biased. To mitigate this, we impute missing view posteriors at the distribution level, ensuring a stable, reliable aggregated posterior that better captures the data distribution.
@@
We first compute the aggregated posterior from available views using Eq. (14), then identify its $k$ nearest neighbors $\mathcal{K}_{i,v}$ (with view $v$ present) based on the 2-Wasserstein distance between diagonal Gaussian posteriors:
@@
The clustering is performed by computing posterior responsibilities using reparameterization trick:
@@
After selecting feasible positions for imputation, the variational inference framework further estimates the uncertainty of the imputation process. This uncertainty-aware strategy dynamically adjusts the contributions of different views during fusion, mitigating the influence of low-quality observations and unreliable imputations on the final representation.
III-E Training Objective.
We adopt a variational training objective that combines the standard Evidence Lower Bound (ELBO) [kingma2013auto] with a cross-view coherence regularization to optimize the model. This formulation enables effective multi-view representation learning while handling missing views and preserving latent structure.
@@
1:Input: Incomplete multi-view dataset $\{\{\bm{x}_{i}^{v}\}_{v=1}^{V}\}_{i=1}^{N}$ with indicator matrix $\mathbf{M}$; Number of clusters $K$; Information score threshold $\tau$; Regularization parameter $\alpha$.
@@ IV-A Experimental Setup
We evaluate our method on four widely used real-world multi-view datasets: (1) Caltech7-5V [fei2004learning, li2015large] contains 1474 pictures of objects belonging to 7 classes and all images are described with five types of features (254 CENTRIST, 512 GIST, 1984 HOG, 928 LBP, and 40 wavelet moments). (2) Scene-15 [fei2005bayesian] consists of 4,485 samples from 15 commonly seen indoor and outdoor scene categories, with each category containing 200–400 images features. It provides diverse scene samples covering natural environments (e.g., coast, forest), man-made environments (e.g., office, bedroom), and mixed scenes (e.g., street, store). (3) COIL100 [nene1996columbia] includes 7,200 images from 100 object categories, and each object is captured from different angles. Three types of features are used as distinct views, namely Isometric Projection (ISO), Linear Discriminant Analysis (LDA), and Neighborhood Preserving Embedding (NPE). (4) Multi-Fashion [xiao2017fashion] is a multi-view extension of the Fashion-MNIST dataset, where different feature extractors are used to construct heterogeneous representations for each image. The details of these datasets are summarized in TABLE I.
@@
Dataset Instance Cluster Dimensions
@@
We compare our approach with seven IMC baselines, which can be categorized into three main types based on their strategies for handling missing data: imputation-based methods, imputation-free methods, and cautious imputation methods:
  • Imputation based methods: PMIMC [yuanPrototype2025], the current state-of-the-art among imputation-based methods, performs prototype-based imputation by aligning incomplete samples with learned prototypes across views. CPSPAN [jinDeep2023] employs a structure embedding imputation strategy to align features between views by filling in missing embeddings using nearby neighbors from other views.
  • Imputation-free methods: GIMVC [baiGraphguided2024] maximizes the utilization of existing information by fusing feature and graph information from observed samples. DIMVC [xu2022deep] avoids both imputation and feature fusion, instead projecting multi-view representations into a shared high-dimensional space through nonlinear mappings. DVIMC [xu2024deep], often achieving competitive or superior performance among imputation-free approaches, aggregates representations from observed views using the Product-of-Experts (PoE) approach.
  • Cautious imputation methods: DCP [linDual2022] performs contrastive learning only on complete samples and applies imputation before evaluation. DSIMVC [tangDeepSafeIncomplete2022] learns a weighting function that down-weights unreliable imputed samples to mitigate the negative impact of inaccurate imputations.
@@
To better reflect real-world conditions, this work investigates incomplete multi-view clustering under unbalanced and more challenging missing scenarios. Instead of setting a global ratio of incomplete instances, we simulate such situations by randomly removing views from each instance with different probabilities, and define the missing rate as the proportion of missing positions across all views, which is more standard and stricter than instance-level ratio. Clustering performance is assessed with Accuracy (ACC), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI), averaged over ten runs. For hyperparameters, we empirically set the regularization coefficient $\alpha$ to 5 for Caltech7-5V and COIL100, 10 for Multi-Fashion and Scene-15. All experiments run on an NVIDIA 5070Ti GPU.
@@ IV-B Experimental Evaluation
To comprehensively evaluate the effectiveness, interpretability, and generality of the proposed Information-Guided Selective Multi-View Clustering (ISMVC) framework, we design experiments to answer the following key questions:
  • Q1: Overall Effectiveness. Does ISMVC achieve superior clustering performance compared to state-of-the-art imputation-based and imputation-free methods under various missing rates?
  • Q2: Informativeness and Interpretability. How do information scores reflect the heterogeneity of missing positions, and how effectively do they guide the selective imputation process?
  • Q3: Robustness and Imputation Ratio Sensitivity. How robust is ISMVC to different degrees of missingness and varying selective imputation ratios? Does the new imputation mechanism mitigate performance degradation under high missing rates?
  • Q4: Generality and Plug-in Capability of IBSI. Can the proposed IBSI module be seamlessly integrated into other IMVC frameworks to enhance their robustness and accuracy?
  • Q5: Sensitivity to the Regularization Coefficient. How does the regularization parameter $\alpha$ affect the clustering performance of ISMVC?
@@
  • Robustness under severe missingness. While imputation-free methods (e.g., DVIMC) exhibit drastic degradation when missing rates are high (e.g., ACC on Caltech7-5V drops from 0.872 to 0.516 as the missing rate increases from 0.1 to 0.5), our method maintains stable performance, demonstrating that the proposed informativeness-based imputation position selection strategy is especially valuable when available evidence becomes sparse. Detailed visualized analysis follows.
@@
(a) Heatmap of position-level information scores
@@
Figure 5: Visualization of information score on Scene-15. (a) Heatmap of position-level scores with $\eta=0.5$, where deeper blue indicates higher informativeness. The darkest cells with diagonal hatching correspond to observed entries, which are excluded from computation. (b) Violin plots of information score distributions across different missing rates. As missing rate increases, the score distribution shifts toward lower values and becomes more polarized.
@@
To better understand how information scores guide selective imputation, we provide both position-level and distribution-level visualizations on the Scene-15 dataset (Fig. 5). These analyses jointly reveal the heterogeneous informativeness of missing positions and its evolution with varying missing rates, offering direct evidence for our design motivation.
@@
Fig. 5a presents the heatmap of information scores under a 0.5 missing rate, where each cell denotes the score of a missing position. Darker shades correspond to higher score, while the darkest cells with diagonal hatching indicate observed data that do not require imputation. The missing pattern is configured such that view 1 suffers from the highest probability of missingness, followed by view 2 and view 3. This visualization provides a fine-grained perspective on how the availability of intra-view neighbors and cross-view correspondences influences the informativeness of each missing position.
@@
In detail, the heatmap demonstrates clear heterogeneity across views: missing positions in view 2 or view 3 often achieve higher scores (e.g., positions (138, 2) and (801, 3)), owing to the availability of more intra-view neighbors and stronger cross-view support. In contrast, samples missing in view 1 typically exhibit lower informativeness, as the heavy missingness reduces the size of the support set. These observations suggest that samples missing in view 1 with limited cross-view evidence are inherently less suitable for imputation, highlighting the necessity of an Informativeness-based Imputation position selection mechanism rather than indiscriminate filling.
@@
Beyond individual cases, we further examine how information scores are distributed across different missing rates. Figure 5b shows violin plots of the score distributions. To account for the increasing number of missing positions at higher missing rates, the distributions are plotted based on proportions rather than absolute counts, facilitating a fair comparison across varying levels of incompleteness.
@@
Several trends are noteworthy in the violin plots. First, as the missing rate increases from 0.1 to 0.5, the mean information score drops sharply (from approximately 4974 to 1108), and the overall distribution shifts downward, reflecting the diminishing quantity and reliability of available support. Second, at low missing rates ($\eta$ = 0.1, 0.2), the scores concentrate into a single sharp peak at high values, indicating that most missing positions can be confidently estimated. As the missingness grows ($\eta$ = 0.3, 0.4, 0.5), the distribution gradually flattens and splits into two modes—one corresponding to well-supported positions and the other to poorly-supported ones—with the lower mode becoming dominant at $\eta$ = 0.5. This transition from a unimodal to a pronounced bimodal pattern reveals the increasing heterogeneity in imputation reliability and highlights the necessity of selectively imputing positions with sufficiently high information scores.
@@
The above findings provide strong empirical justification for our method. The position-level heatmap demonstrates that information availability is highly uneven across different views and positions, while the violin plots reveal that such unevenness grows into polarization as missingness increases. Together, these analyses confirm that imputation should not be uniformly applied. Instead, it is crucial to prioritize positions with abundant and reliable evidence while avoiding those in the low-score regime.
@@
To further examine the effect of selective imputation, we study how varying the proportion of imputed positions influences clustering performance. Specifically, we adjust the selection ratio by changing the threshold of information scores, and conduct experiments on Caltech7-5V and Scene-15 under both low ($\eta=0.1$) and high ($\eta=0.5$) missing rates. The results are presented in Fig. 6.
@@
In this section, we further evaluate the generality and plug-in potential of the proposed Informativeness-Based Selective Imputation (IBSI) by integrating it into two representative frameworks with distinct philosophies: the imputation-free GIMVC and the complete-sample-only DCP. For fairness, both baselines use the same simple cross-view neighbor mean strategy in the raw feature space, ensuring that IBSI is the only additional component.
@@
A further observation is that conservative selection ratios—i.e., only imputing positions with sufficiently high informativeness—yield significantly better performance than aggressive imputations. This reflects the risk of introducing noise through unreliable estimates and highlights the benefit of focusing on high-confidence entries. Owing to the simplicity of the plug-in imputation mechanism, the optimal ratio is around 0.3, in contrast to about 0.5 in ISMVC where the imputation process is more principled and distribution-aware. This suggests that the appropriate selection ratio should be adapted to the sophistication of the underlying imputation module: simple schemes favor conservative ratios, while more advanced models can safely leverage a higher proportion of missing entries.

### v2

@@ III-B1 The Multi-view Generative Assumption
In the generative process described above, the cluster assignment $\bm{c}_{i}$ first determines the latent representation $\bm{z}_{i}$, which serves as the underlying variable that generates multi-view observations $\bm{x}_{i}^{v}$}. During inference, the direction is reversed: we aim to infer $\bm{z}_{i}$ and $\bm{c}_{i}$ from the observed views $\bm{x}_{i}^{v}$}. Specifically, we first infer a latent representation $\bm{z}_{i}$ by integrating information from all available views, and then estimate the cluster assignment $\bm{c}_{i}$ based on the $\bm{z}_{i}$. Hence, the learning objective focuses on obtaining accurate and stable latent representations rather than reconstructing all missing observations. However, incomplete views provide limited evidence, which leads to uncertain or biased estimates of posterior distribution of $\bm{z}_{i}$ and consequently degrades clustering performance.
There are two intuitive strategies to handle missing data. One is to directly recover missing data from the latent representation $\bm{z}_{i}$. However, this imputation paradigm contributes little to learning a more discriminative and cluster-consistent $\bm{z}_{i}$, and thus brings limited benefit to our clustering objective. Another strategy is to ignore the missing views and only aggregate observed ones [39]. Yet under high missing rates or severe imbalance, this approach suffers from drastic degradation, and in extreme cases degenerates into a single-view solution that still outputs over-confident but biased representations.
@@
To address this, we introduce two complementary components: pre-imputation assessment of missing positions and variational inference with selective imputation. The former identifies positions with abundant and trustworthy evidence, ensuring that only well-supported missing positions are imputed, thereby mitigating the negative impact of unreliable auxiliary information. The latter models the uncertainty inherent in the imputation process by aggregating information from local neighbors at the distributional level, which not only stabilizes the missing-view likelihood approximation but also facilitates the learning of more robust and cluster-consistent latent representations.
III-B2 Informativeness-Based Imputation Position Selection
@@
Building upon these two complementary evidence sources, we construct a unified estimation framework that quantifies their information contribution via a position-specific support set. This allows us to compute an informativeness score $\mathrm{Info}(i,v)$ for each missing position and perform imputation only when sufficient support is available.
This quantification is inspired by the generative assumption, where samples with more similar latent posteriors are more likely to belong to the same Gaussian component (i.e., cluster). From this perspective, the contribution of different reference samples and views to imputing a missing position $\tilde{\bm{x}}_{i}^{v}$ can be naturally distinguished:
@@
We quantify the imputation-relevant informativeness of missing positions via the above principles. Each sample-view pair contributes differently based on its availability, similarity, and alignment with the target view. Specifically, the informativeness-based imputation position selection includes defining support samples, estimating information contributions, and computing informativeness scores. Next, each part will be introduced in detail.
@@
In practice, not all samples are suitable for participating in the quantification of imputation-relevant informativeness for a missing position. As illustrated on the right side of Fig. 3, samples that exhibit completely different missing patterns or lack any overlap in observed views should be excluded. This exclusion is not due to unreliability of the data, but rather because these samples do not provide a clear pathway for estimating the target view.
In most imputation paradigms, the estimation of a missing position relies on observable correspondences: either intra-view neighbors that share similar features, or cross-view relations established through jointly observed samples. When two samples share no common observed view, their potential statistical dependencies can only be implicitly captured during global network training, which is indirect, unstable, and not interpretable at the sample level. Therefore, we construct a set of support samples that serve as valid reference points for each missing position $(i,v)$.
@@
1. View correlation. To quantify the consistency of shared semantics across views and provide reliable weighting for similarity approximation when the target view is missing, the inter-view correlation $\mathrm{corr}^{uv}\in(0,1]$ is computed in the latent space via Canonical Correlation Analysis (CCA) [1]. Specifically, for views $u$ and $v$, we consider the $N_{uv}=|\mathcal{X}^{u}\cap\mathcal{X}^{v}|$ samples observed in both views and form their latent representations:
@@
where $\bm{\mu}^{v}$ denotes the latent embedding output by the encoder $f^{v}(\bm{x}^{v})$. At this stage, $\bm{\mu}^{v}$ is a deterministic representation without probabilistic semantics, and the pre-training is solely used to obtain a well-initialized latent space for reliable imputation-relevant informativeness estimation and efficient optimization. The same encoder–decoder architecture is later extended to a variational framework for uncertainty-aware modeling. Note that while using CCA, our framework is agnostic to the specific choice of correlation measure. For different models, better options such as mutual information or cross-view mapping loss could be preferred.
2. Sample similarity. To differentiate the contributions of intra-view neighbors and prioritize trustworthy samples more likely to belong to the same cluster or Gaussian component, intra-view similarity is defined as a normalized distance metric. Specifically, intra-view similarity $\mathrm{sim}_{ij}^{u}$ is computed as:
@@
Informativeness Score Computation
@@
For cross-view evidence $\bm{x}_{i}^{u}$ ($u\neq v$), the effectiveness of these evidence hinges on the quality of view-to-view semantic transfer, such as how well the mapping between views $u$ and $v$ has been learned. As this direct quality is challenging to measure, we approximate the contribution of $\bm{x}_{i}^{u}$ indirectly by aggregating Informativeness score from co-occurring support samples weighted by view correlation and local similarity:
@@
By setting $\mathrm{corr}^{vv}=1$, the total Informativeness score for $\tilde{\bm{x}}_{i}^{v}$ is the sum of intra- and cross-view score:
@@
We use this scalar value to selectively impute only those missing positions that are sufficiently supported. In detail, we assign the threshold $\tau$ connected with a proportion of missing data, and only perform imputation when $\mathrm{Info}(i,v)>\tau$.
III-B3 Variational Inference with Selective Imputation
@@
Following the mean-field factorization, we first focus on learning a shared latent representation $\bm{z}_{i}$, which serves as the basis for subsequent clustering.
Common Representation Learning
@@
where $\bm{\mu}_{i}$ and $(\bm{\sigma}_{i})^{2}$ denote the aggregated mean and variance across available views $\mathcal{V}_{i}$. This weights views by confidence (lower variance), but high missing rates or unbalanced view missing can destabilize the posterior. Missing views reduce terms in Eq. (15), skewing $\bm{\mu}_{i}$ and underestimating $(\bm{\sigma}_{i})^{2}$, especially when views are systematically absent, leading to biased estimates. To mitigate this, we impute missing view posteriors at the distribution level, ensuring a stable, reliable aggregated posterior that better captures the data distribution.
@@
We first compute the parameters of aggregated posterior from available views using Eq. (15), then identify its $k$ nearest neighbors $\mathcal{K}_{i,v}$ (with view $v$ present) based on the 2-Wasserstein distance between diagonal Gaussian posteriors:
@@
Based on the stabilized latent posterior $q(\bm{z}_{i})$, we further infer the shared cluster assignment. Specifically, a latent sample $\bm{z}_{i}^{(1)}$ is drawn via the reparameterization trick, and the posterior responsibilities are then computed accordingly:
@@
After identifying feasible positions for imputation, the variational inference framework is employed to explicitly model the uncertainty associated with the imputation process. This uncertainty-aware mechanism adaptively modulates the contributions of different views during fusion, thereby suppressing the impact of low-quality observations and unreliable imputations on the learned representations.
III-C Loss Function
We adopt a variational training objective that combines the standard Evidence Lower Bound (ELBO) [17] with a cross-view coherence regularization to optimize the model. This formulation enables effective multi-view representation learning while handling missing views and preserving latent structure.
@@ III-C Loss Function
1:Input: Incomplete multi-view dataset $\{\{\bm{x}_{i}^{v}\}_{v=1}^{V}\}_{i=1}^{N}$ with indicator matrix $\mathbf{M}$; Number of clusters $K$;Informativeness score threshold $\tau$; Regularization parameter $\alpha$.
@@ IV-A Experimental Setup
We evaluate our method on four widely used real-world multi-view datasets: (1) Caltech7-5V [9, 19] contains 1474 pictures of objects belonging to 7 classes and all images are described with five types of features (254 CENTRIST, 512 GIST, 1984 HOG, 928 LBP, and 40 wavelet moments). (2) Scene-15 [10] consists of 4,485 samples from 15 commonly seen indoor and outdoor scene categories, with each category containing 200–400 images features. It provides diverse scene samples covering natural environments (e.g., coast, forest), man-made environments (e.g., office, bedroom), and mixed scenes (e.g., street, store). (3) COIL100 [26] includes 7,200 images from 100 object categories, and each object is captured from different angles. Three types of features are used as distinct views, namely Isometric Projection (ISO), Linear Discriminant Analysis (LDA), and Neighborhood Preserving Embedding (NPE). (4) Multi-Fashion [37] is a multi-view extension of the Fashion-MNIST dataset, where different feature extractors are used to construct heterogeneous representations for each image. The details of these datasets are summarized in TABLE I.
@@
Dataset Sample Cluster Dimensions
@@
We compare our approach with seven IMC baselines, which can be categorized into three main types based on their strategies for handling missing data: imputation-based methods, imputation-free methods, and selective imputation methods:
  • Imputation based methods: PMIMC [48], the current state-of-the-art among imputation-based methods, performs prototype-based imputation by aligning incomplete samples with learned prototypes across views. CPSPAN [16] employs a structure embedding imputation strategy to align features between views by filling in missing embeddings using nearby neighbors from other views.
  • Imputation-free methods: GIMVC [3] maximizes the utilization of existing information by fusing feature and graph information from observed samples. DIMVC [41] avoids both imputation and feature fusion, instead projecting multi-view representations into a shared high-dimensional space through nonlinear mappings. DVIMC [39], often achieving competitive or superior performance among imputation-free approaches, aggregates representations from observed views using the Product-of-Experts (PoE) approach.
  • Selective imputation methods: DCP [20] performs contrastive learning only on complete samples and applies imputation before evaluation. DSIMVC [31] learns a weighting function that down-weights unreliable imputed samples to mitigate the negative impact of inaccurate imputations.
@@
To better reflect real-world conditions, this work investigates incomplete multi-view clustering under unbalanced and more challenging missing scenarios. Instead of setting a global ratio of incomplete samples, we simulate such situations by randomly removing views from each sample with different probabilities, and define the missing rate as the proportion of missing positions across all views, which is more standard and stricter than sample-level ratio. Clustering performance is assessed with Accuracy (ACC), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI), averaged over ten runs. For hyperparameters, we empirically set the regularization coefficient $\alpha$ to 5 for Caltech7-5V and COIL100, 10 for Multi-Fashion and Scene-15. All experiments run on an NVIDIA 5070Ti GPU.
@@ IV-B Experimental Evaluation
To comprehensively evaluate the effectiveness, interpretability, and generality of the proposed Implicit Informativeness-based Selective Imputation (SI3) method, we design experiments to answer the following key questions:
  • Q1: Overall Effectiveness. Does SI3 achieve superior clustering performance compared to state-of-the-art imputation-based and imputation-free methods under various missing rates?
  • Q2: Informativeness and Interpretability. How do Informativeness scores reflect the heterogeneity of missing positions, and how effectively do they guide the selective imputation process?
  • Q3: Robustness and Imputation Ratio Sensitivity. How robust is SI3 to different degrees of missingness and varying selective imputation ratios? Does the new imputation mechanism mitigate performance degradation under high missing rates?
  • Q4: Generality and Plug-in Capability of IBSI. Can the proposed IBSI module be seamlessly integrated into other IMC frameworks to enhance their robustness and accuracy?
  • Q5: Sensitivity to the Regularization Coefficient. How does the regularization parameter $\alpha$ affect the clustering performance of SI3?
@@
  • Robustness under severe missingness. While imputation-free methods (e.g., DVIMC) exhibit drastic degradation when missing rates are high (e.g., ACC on Caltech7-5V drops from 0.872 to 0.516 as the missing rate increases from 0.1 to 0.5), our method maintains stable performance, demonstrating that the proposed informativeness-based imputation position selection strategy is especially valuable when available evidence becomes sparse. Detailed visualized analysis follows in Q2.
@@
(a) Heatmap of position-level informativeness scores
@@
Figure 5: Visualization of informativeness score on Scene-15. (a) Heatmap of position-level scores with $\eta=0.5$, where deeper blue indicates higher score. Cells with diagonal hatching correspond to observed position, which are excluded from computation. (b) Violin plots of informativeness score distributions across different missing rates. As missing rate increases, the score distribution shifts toward lower values and becomes more polarized.
@@
To better understand how informativeness score guide selective imputation, we provide both position-level and distribution-level visualizations on the Scene-15 dataset (Fig. 5). These analyses jointly reveal the heterogeneous informativeness of missing positions and its evolution with varying missing rates, offering direct evidence for our design motivation.
@@
Fig. 5a presents the heatmap of informativeness scores under a 0.5 missing rate, where each cell denotes the score of a missing position. Darker shades correspond to higher score, while the darkest cells with diagonal hatching indicate observed data that do not require imputation. The missing pattern is configured such that view 1 suffers from the highest probability of missingness, followed by view 2 and view 3. This visualization provides a fine-grained perspective on how the availability of intra-view neighbors and cross-view correspondences influences the imputation-relevant informativeness of each missing position.
@@
In detail, the heatmap demonstrates clear heterogeneity across views: missing positions in view 2 or view 3 often achieve higher scores (e.g., positions (138, 2) and (801, 3)), owing to the availability of more intra-view neighbors and stronger cross-view support. In contrast, samples missing in view 1 typically exhibit lower informativeness score, as the heavy missingness reduces the size of the support set. These observations suggest that samples missing in view 1 with limited cross-view evidence are inherently less suitable for imputation, highlighting the necessity of an Informativeness-based Imputation position selection mechanism rather than indiscriminate filling.
@@
Beyond individual cases, we further examine how informativeness scores are distributed across different missing rates. Figure 5b shows violin plots of the score distributions. To account for the increasing number of missing positions at higher missing rates, the distributions are plotted based on proportions rather than absolute counts, facilitating a fair comparison across varying levels of incompleteness.
@@
Several trends are noteworthy in the violin plots. First, as the missing rate increases from 0.1 to 0.5, the mean informativeness score drops sharply (from approximately 4974 to 1108), and the overall distribution shifts downward, reflecting the diminishing quantity and reliability of available support. Second, at low missing rates ($\eta$ = 0.1, 0.2), the scores concentrate into a single sharp peak at high values, indicating that most missing positions can be confidently estimated. As the missingness grows ($\eta$ = 0.3, 0.4, 0.5), the distribution gradually flattens and splits into two modes—one corresponding to well-supported positions and the other to poorly-supported ones—with the lower mode becoming dominant at $\eta$ = 0.5. This transition from a unimodal to a pronounced bimodal pattern reveals the increasing heterogeneity in imputation reliability and highlights the necessity of selectively imputing positions with sufficiently high informativeness score.
@@
The above findings provide strong empirical justification for our method. The position-level heatmap demonstrates that imputation-relevant informativeness is highly uneven across different views and positions, while the violin plots reveal that such unevenness grows into polarization as missingness increases. Together, these analyses confirm that imputation should not be uniformly applied. Instead, it is crucial to prioritize positions with abundant and reliable evidence while avoiding those in the low-score regime.
@@
To further examine the effect of selective imputation, we study how varying the proportion of imputed positions influences clustering performance. Specifically, we adjust the selection ratio by changing the threshold of informativeness scores, and conduct experiments on Caltech7-5V and Scene-15 under both low ($\eta=0.1$) and high ($\eta=0.5$) missing rates. The results are presented in Fig. 6.
@@
In this section, we further evaluate the generality and plug-in potential of the proposed Informativeness-Based Selection and Indication (IBSI) module by integrating it into two representative frameworks with distinct philosophies: the imputation-free GIMVC and the complete-sample-only DCP. For fairness, both baselines use the same simple cross-view neighbor mean strategy in the raw feature space, ensuring that IBSI is the only additional component.
@@
A further observation is that conservative selection ratios—i.e., only imputing positions with sufficiently high informativeness score—yield significantly better performance than aggressive imputations. This reflects the risk of introducing noise through unreliable estimates and highlights the benefit of focusing on high-confidence positions. Owing to the simplicity of the plug-in imputation mechanism, the optimal ratio is around 0.3, in contrast to about 0.5 in SI3 where the imputation process is more principled and distribution-aware. This suggests that the appropriate selection ratio should be adapted to the sophistication of the underlying imputation module: simple schemes favor conservative ratios, while more advanced models can safely leverage a higher proportion of missing positions.